The equilibrium constant for the reaction Sr(s) + Mg2+(aq) ⇌ Sr2+(aq) + Mg(s) is 3.69 × 1017 at 25°C. Calculate E o for a cell made up of Sr / Sr2+ and Mg / Mg2+ half-cells. V
Answer #1


use below formula for finding Eo from K
Eo = (2.303*R*T)/(n*F) log Kc
At 25 oC or 298 K, 2.303*R*T/F = 0.0592
So, Eo = (0.0592/n)*log Kc
Eo = (0.0592/2)*log(3.69*10^17)
Eo = 0.520 V
Answer: 0.520 V